OpenStudy (anonymous):

Sandy is working with a carpenter to frame a house. They are using 8 foot long boards but each board must be cut to be 7 feet, 10 3/4 inches long. How much is cut off each board?

OpenStudy (anonymous):

12 - 10 3/4 = 1 1/4 inches = 5/4 inches

OpenStudy (anonymous):

thanks {:

OpenStudy (anonymous):

YW!!:) And Welcome to Open Study!!! Let me know if you need any more help with anything!:)